A car trunk lies on a horizontal rough floor. The coefficient of friction between the truck and the floor is 0.4330.calculate the magnitude of the force p which is necessary to pull the truck horizontally if the force is applied a. Horizontally b.At 30°above the horizontal
1
Expert's answer
2019-11-07T09:07:53-0500

Given:

μ = 0.433

α = 30°

m


Find: Pa, Pb


a)

"f_{k_x} + mg_x + N_x + P_x = 0\\
P = \mu mg = 0.433\ *\ 10m/s^2 \ *\ m\ kg =\ 4.33m\ N"


b)

"f_{k_x} + mg_x + N_x + P_x = 0\\
\mu mg = Pcos\alpha\\
P = \dfrac{\mu mg}{cos\alpha} = \dfrac{0.433 * 10m/s^2\ *\ m\ kg}{cos 30\degree} \approx 5m\ N"

Answer: a) Pa=4.33m N b) Pb=5m N

(where m is the mass of a truck in kilos)